OpenStudy (anonymous):

If there are 21 values in a data set in order from smallest to largest, what is the first quartile of the data set? A. The mean of the 3rd and 4th values B. The mean of the 4th and 5th values C. The mean of the 5th and 6th values D. The mean of the 6th and 7th values

OpenStudy (anonymous):

@mathmale @wio

OpenStudy (anonymous):

@heacain

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Well, I would try \(21/4\).

OpenStudy (anonymous):

5.25

OpenStudy (anonymous):

That gives you \(5.25\), which leads me to think it is between \(5\) and \(6\).

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

thanks man u smart

OpenStudy (anonymous):

but where did the 4 come from?

OpenStudy (anonymous):

because i have another problem very similar and still dont know how to do it

OpenStudy (anonymous):

There are 4 quartiles

OpenStudy (anonymous):

If there are 21 values in a data set in order from smallest to largest, what is the third quartile of the data set? A. The mean of the 17th and 18th values B. The mean of the 18th and 19th values C. The mean of the 16th and 17th values D. The mean of the 19th and 20th values

OpenStudy (anonymous):

so i do 3 this time?

OpenStudy (anonymous):

because they wanna know the third quaritle?

OpenStudy (anonymous):

For third quartile, you would use \(\frac 34\)

OpenStudy (anonymous):

First quartile is \(\frac 14\) and second quartile is \(\frac 12\)

OpenStudy (anonymous):

21/34?

OpenStudy (anonymous):

i dont understand im sorry

OpenStudy (anonymous):

Okay, do you understand how to find the median?

OpenStudy (tennis5518):

Im late! :(

OpenStudy (anonymous):

smallest to largest the middle number?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Now when you split things in half, the first quartile is the median of the first half.

OpenStudy (anonymous):

If there are 21 numbers, then the media is at 11. 1-10 is ten numbers, 12-21 is ten numbers.

OpenStudy (anonymous):

im following

OpenStudy (anonymous):

First quartile is the median of 1-10 numbers.

OpenStudy (anonymous):

second is the 12-21

OpenStudy (anonymous):

1-5 is five numbers, 6-10 is five numbers. This means the median is the median of 5 and 6.

OpenStudy (anonymous):

We also split 12-21 the same way. We have 12-16 is five number, and 17-21 is five numbers.

OpenStudy (anonymous):

im lost now

OpenStudy (anonymous):

``` 12 13 14 15 16 17 18 19 20 21 1 2 3 4 5 5 4 3 2 1 ```

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

so it would be C?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how do i find the third quaritle

OpenStudy (anonymous):

median of later half of the list